Recent Log Entries

CREW CHIEF SIGN IN

Log Number: 4126020Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 02-07-23
Time : 07:39:02
Day : Tuesday
Shift : Day

Crew Chief : Carlino
Program Deputy: Letta
Team : X
Arm : RADCON
Designated SSO: Schoene
LERF Operator:
Operators : Schaumloffel, Schoene

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

FSD Masking Tool: Successful System Masking for: BeamLossMonitor

Log Number: 4126012Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: kiarra
Areas: All

BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM1A09
ILM5A10
ILM1A16
ILM1A19
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09

LEM North Linac

Log Number: 4126002Logbook(s): Tags: Body: LEM 9.15 was successfully run for the North Linac on Tue Feb 7 06:53:50 2023. The application phase took 0:00:14 to complete.

The following cavities have been GSET 0: 1L03-2

Log Number: 4125997Logbook(s): Tags: Body: 

The following cavities have been GSET 0:
1L03-2
Username: kiarra
Date/Time: 02-07-23_06:47:45
RF Status: Bad
Tuner Status: Good
Comment: GSET 0 - stepped off +100k - PMES all over the place, had issues during RF recovery
Special Instructions:

something is changing Hall A beam without intervention

Log Number: 4125991Logbook(s): Tags: Entry Makers: kpriceBody: 

current to Hall A is slowly draining off as I try to tune their optics - then there's a discharge in the injector, positions are thrown into disarray, and the beam profile changes. It's becoming almost impossible to tell which profile changes are due to my quad adjustments and which are due to fluctuations elsewhere in the machine.

Fig 1: positions are all over the place, current drains off while sending tune beam.

Fig 2: post-discharge excursion in y position. -0.10 is nominal.

Fig 3: beam sizes for consecutive harp scans with no quad changes.

Still frames from video showing 1A SLM size oscillation

Log Number: 4125981Logbook(s): Entry Makers: eformanBody: 

Still frames from video showing 1A SLM size oscillation. The green box superimposed on the image is the beam size at its smallest point in the oscillation. There was no tune beam in the machine at the time, only CW to halls C & D.

beam loss excursions grew overnight

Log Number: 4125980Logbook(s): Entry Makers: beneschBody: 

It is interesting that the BLA loss signal, blue in middle pane, grew overnight with fixed currents A at 5 uA and C at 40 uA. I don't know what this means. It remains reasonable symmetric. No loss evident on NDX detectors, top and bottom panes. Ten BLA trips on owl shift, figure 2.

Pages